OpenStudy (anonymous):

A place-kicker must kick a football from a point 36.0 m (about 40 yards) from the goal. Half the crowd hopes the ball will clear the crossbar, which is 3.05 m high. When kicked, the ball leaves the ground with a speed of 24.0 m/s at an angle of 46.0° to the horizontal. (a) By how much does the ball clear or fall short (vertically) of clearing the crossbar? (Enter a negative answer if it falls short.)

OpenStudy (dumbcow):

\[x(t) = 24\cos(46) t\] \[y(t) = -4.9t^2 +24\sin(46)t\] setting y=0 will tell you how long football is in air t = 3.52 plug thatinto x(t) to find how far the football goes x = 58.68 so it goes way past the required 36m now see if it clears the goalpost find time when x = 36, plug that into y(t) t=2.15 y = 14.4 so it clears by about 11.35m